log16(4)-log27(x)=(1/6) solve for x
Answer by stanbon(75887)

log16(4)-log27(x)=(1/6) solve for x
----
[log(4)/log(16)] - log27(x) = 1/6
----------------------------
2 - log27(x) = 1/6
---------
log27(x) = 11/6
---
x = 27^(11/6)
----
x = 425.89
-------------
Cheers,
Stan H.
----------------